# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v1

@@
Abstract
Decomposing visual scenes into objects, as humans do, facilitates modeling object relations and dynamics. Object-Centric Learning (OCL) achieves this by aggregating image or video feature maps into object-level feature vectors, known as slots. OCL’s self-supervision via reconstructing the input from slots struggles with complex textures, thus many methods employ Vision Foundation Models (VFMs) to extract feature maps with better objectness. However, using VFMs merely as feature extractors does not fully unlock their potential. We propose Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO), where VFM features are extracted to facilitate object-level information aggregation and further quantized to strengthen supervision in reconstruction. Our VVO unifies OCL representatives into a concise architecture. Experiments demonstrate that VVO not only outperforms mainstream methods on object discovery tasks but also benefits downstream tasks like visual prediction and reasoning. The source code is available in the supplement.
Object-Centric Learning, Vision Foundation Models, Vector Quantization, Object Discovery, Visual Prediction, Visual Reasoning
Aalto University, Espoo, Finland
{rongzhen.zhao, vivienne.wang, juho.kannala, joni.pajarinen}@aalto.fi
@@ Abstract
Figure 1: We utilize two observations: (i) Compared to non-VFMs, VFMs extract features with better object separability, i.e., smaller intra-object distances and larger inter-object distances $\Rightarrow$ We facilitate OCL aggregation via VFM features; (ii) VFM and non-VFM features have a distribution gap, i.e., far-flung centroids $\Rightarrow$ We strengthen OCL supervision by sharing the aforementioned VFM features and quantizing them as the reconstruction target.
1 Introduction
@@
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri & Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a).
@@
Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020). Firstly, the encoder transforms input image or video frame pixels into a dense feature map. Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots. Lastly, the decoder reconstructs the input in some form from these slots, to provide the self-supervised training signal.
@@
Due to its self-supervision nature, OCL heavily relies on pixel textures to separate different objects. The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects with complex textures. Some methods (Kipf et al., 2022; Elsayed et al., 2022) reconstruct the optical flow or depth map of the input to mitigate textural noises, albeit at the cost of extra expensive annotations. Some others (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
However, these methods merely treat VFMs as stronger feature extractors for OCL encoding, overlooking their potential to provide better reconstruction targets for OCL decoding. VFMs have yet to be fully utilized in OCL.
We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, VQ-VFM-OCL or VVO. As shown in Figure 1, we fully utilize VFMs in OCL in two ways. We not only use VFMs to extract feature maps with better objectness, to facilitate object information aggregation. We also quantize VFM feature maps with codebook codes, to strengthen supervision in reconstruction. Such shared quantization suppresses redundant feature information and produces consistent reconstruction targets across samples, facilitating aggregation via those initial slots, which are also shared across samples.
Our contributions are: (i) A unified architecture of encoder-aggregator/quantizer-decoder, which supports various OCL modules; (ii) Three key designs of direct VFM feature extraction, shared VFM feature quantization to fully utilize VFMs in OCL, and a VQ variant designed for OCL; (iii) Significant improvements in mainstream OCL methods, along with benefits to downstream task performance.
2 Related Work
@@
SSseg vs OCL. Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024) and OCL can discovery objects, i.e., segment object, in images or videos without supervision. But SSseg only outputs object segmentation masks, while OCL represents objects as slots with segmentation as pyproduct. Besides, these slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks, like prediction, reasoning, planning and decision-making. A direct comparison between OCL and SSseg methods is beyond the scope of this work.
@@
OCL encoding. The stronger encoder that extracts feature maps of better objectness contributes more to OCL performance. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them in OCL. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also recognized VFMs, they remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
@@
OCL aggregation. SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods. Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are all designed without changing the external interface. But considering their performance boosts, where some are very limited, we only integrate BO-QSA by default.
@@
OCL decoding. With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision. Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (Zadaianchuk et al., 2024), decodes each slot’s spatial broadcast (Watters et al., 2019) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction. Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition. Diffusion-based decoding, used in LSD (Jiang et al., 2023) and SlotDiffusion, drives slots to recover noise added to the input’s VAE representation. Our VVO supports all these types of OCL decoding.
@@
VAE for OCL. Variational Autoencoders (VAEs), like dVAE (Im Im et al., 2017) in SLATE and VQ-VAE (Van Den Oord et al., 2017) in SlotDiffusion, are employed to produce reconstruction targets for OCL training. Since these VAEs are designed for image generation, some methods adapt them for OCL. Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better. MSF (Zhao et al., 2024b) exploits the multi-scale idea from detection and segmentation in OCL with novel VAE designs. Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
@@
Figure 2: VVO is a unified architecture that fully utilizes VFMs in OCL. It not only extracts VFM features with better objectness to facilitate object information aggregation; but further quantizes those VFM features as reconstruction targets to strengthen OCL training supervision. With typical SlotAttention or it variants as the aggregator and Vector-Quantization as the quantizer, VVO supports different VFMs as the encoder, and supports mainstream mixture, auto-regression and diffusion models as the decoder.
3 Proposed Method
@@
We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, or VQ-VFM-OCL (VVO), unifying mainstream OCL with significant performance boosts. Using VFMs for better performance is not new, but VVO is the first to fully utilize VFMs in the OCL setting.
3.1 Unify OCL
@@
$\bm{\phi}_{\mathrm{d}}:(\bm{S},\bm{Q})\rightarrow\bm{Q}^{\prime}$ (4)
where $\bm{\phi}_{\mathrm{d}}$ can be parameterized as a CNN or MLP if using mixture decoding for OCL (Kipf et al., 2022; Seitzer et al., 2023), where each of $\bm{S}$ is spatially broadcast (Watters et al., 2019) into the shape of $\bm{Q}$ and then decoded into a component and finally mixed together with predicted weights; or (ii) a Transformer decoder if using auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where causally masked $\bm{Q}$ is the query and $\bm{S}$ is the key and value; or (iii) a conditional Diffusion model if using diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ with noise is the input and $\bm{S}$ is the condition.
Reconstructing the destructed $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible. Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important. We will explain this in Section 3.2.
@@
$\mathrm{min}_{(\bm{\phi}_{\mathrm{a}},\bm{\phi}_{\mathrm{d}})}f_{\mathrm{recon}}(\bm{Q}^{\prime},\bm{Q})$ (5)
@@
3.2 Fully Utilize VFMs in OCL
We fully utilize VFMs in the OCL setting by two key designs: (i) direct VFM feature extraction to facilitate object-level information aggregation, and (ii) shared VFM feature quantization as reconstruction targets to strengthen supervision in OCL training.
Figure 3: Better objectness helps OCL aggregation. Green and orange areas stand for objects $\bm{o}_{1}$ and $\bm{o}_{2}$, where $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary super-pixels and $\bm{s}_{*}$ is $\bm{o}_{2}$’s centroid. But the actual query $\bm{s}\sim\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2})$. In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
@@ 3.2 Fully Utilize VFMs in OCL
We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly. No extra position encoding is needed here like in SA (Locatello et al., 2020) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$. Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly. Thus, We expect the subsequent $\bm{\phi}_{\mathrm{a}}$ to aggregate dense $\bm{Z}$ into sparse slots $\bm{S}$ under queries $\bm{S}_{0}$ better.
Direct use of VFMs for feature extraction contributes to better aggregation. As shown in Figure 1, VFM representations show better objectness or object separability than that of non-VFMs, under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.clust
er.KMeans.html
. OCL aggregation based on SlotAttention and its variants is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). This has already been empirically exploited in previous methods like DINOSAUR (Seitzer et al., 2023) and SlotDiffusion (Wu et al., 2023b), but without any explanation. We provide math analysis below.
Math Analysis: VFM Representation Eases Aggregation
$d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$ (6)
[…]
According to Figure 1, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space. This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space. Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space. $\square$
Figure 4: Qualitative object discovery performance comparison.
@@
$\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$ (10)
$\bm{I}=\mathrm{argmax}_{m}(\bm{P})$ (11)
$\bm{Q}=\mathrm{index}_{m}(\bm{T},\bm{I})$ (12)
@@
For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ pre-Equation 10 to encourage code utilization:
$\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$ (13)
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature. We devise annealing residual connection (Zhao et al., 2024a, b) post-Equation 12 to stabilize pre-training:
$\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$ (14)
@@
$l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$ (15)
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency and suppresses information redundancy or noise. This helps the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. After such tokenization, our reconstruction target can be either regressed using mixture and diffusion decoding or classified using auto-regressive decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen supervision.
Math Analysis: Shared VFM Representation Quantization Reduces Supervision Noise
$\frac{\partial\mathrm{MSE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=2(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))\frac{\partial\bm{Q}^{\prime}}{\partial\bm{Z}}$ (16)
$\mathbb{E}[\bm{Q}^{\prime}]=\mathbb{E}[f(\bm{Z})]\approx\bm{Q}$ (17)
$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q})]\approx 0$ (18)
But if instead of sharing $\bm{\phi}_{\mathrm{e}}$, we use an extra VAE encoder plus $\bm{\phi}_{\mathrm{a}}$ to obtain the target, denoted as $\bm{Q}_{2}$, then $\bm{Q}_{2}\neq\bm{Q}$. Substitute $\bm{Q}_{2}$ into Equation 16 and the residual error $\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})$ would be systematically biased:
$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})]\neq 0$ (19)
which increases noise in optimization.
Using CE loss is slightly different but it does not alter our judgment above. $\square$
resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
SLATE-DINO 15.8
$\pm$0.3
85.8
$\pm$0.3

### v3

@@
Abstract.
Perceiving visual scenes as objects and background—like humans do—Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots. OCL’s self-supervision of reconstructing the input from these aggregated slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target. However, existing methods leverage VFM representations in diverse ways and often fail to fully exploit their potential. In response, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods. The key to our unification is simple yet effective, just shared quantizing the same VFM representation as the reconstruction target. Through mathematical modeling and statistical verification, we further analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision. Experiments show that across different VFMs, aggregators and decoders, our VVO consistently outperforms baselines in object discovery and recognition, as well as downstream visual prediction and reasoning. The implementation and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
isbn: 978-1-4503-XXXX-X/2018/06submissionid: 7307copyright: none
@@ Abstract.
Figure 1. We utilize two observations: (i) Compared to non-VFMs, VFMs extract features with better object separability, i.e., smaller intra-object distances and larger inter-object distances $\Rightarrow$ We facilitate OCL aggregation via VFM features; (ii) VFM and non-VFM features have a distribution gap, i.e., separated centroids $\Rightarrow$ We strengthen OCL supervision by reconstructing the quantized features shared from the same VFM rather than from another encoder.
1. Introduction
@@
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024b).
@@
Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020). Firstly, the encoder transforms input image or video frame pixels into a dense feature map. Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots as the query. Lastly, the decoder reconstructs the input in some form from these aggregated slots, to provide the self-supervised training signal.
@@
OCL relies on pixel textures to discover objects. The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects. Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, but at the cost of expensive annotations. Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
However, existing OCL methods leverage VFM representations in quite different ways, as shown in Figure 3, and none of them fully utilize the power of VFM representations.
To address these issues, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods. In this architecture, as shown in Figure 2, our VVO supports different VFMs for encoding, different OCL aggregators and different OCL decoders. The key to such unification is very simple—We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target. We also mathematically and statistically analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
Our contributions are: (i) A clean architecture, which unifies mainstream OCL methods. (ii) Shared quantized VFM representations as reconstruction targets, which not only supports various OCL decoders but also boosts performance. (iii) Insights of why VFM features facilitate OCL aggregation and their shared quantization as reconstruction targets strengthens OCL supervision.
2. Related Work
@@
SSseg vs OCL. Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024a) and OCL can discovery objects, i.e., segmenting objects and the background in images or video frames without supervision. SSseg only outputs object segmentation masks, while OCL further represents objects as slots. These slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks. Thus a direct comparison between them is beyond the scope of this work.
@@
OCL encoding. The stronger encoder that extracts feature maps of better objectness contributes more to OCL performance. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs, they remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
@@
OCL aggregation. SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods. Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are designed without changing the external interface. But considering their performance boosts, we only integrate BO-QSA by default.
@@
OCL decoding. With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision. Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (Zadaianchuk et al., 2024), decodes each slot’s spatial broadcast (Watters et al., 2019) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction. Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition. Diffusion-based decoding in LSD (Jiang et al., 2023) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation. Our VVO supports all these types of OCL decoding.
@@
VAE for OCL. Variational Autoencoders (VAEs), like dVAE (Im Im et al., 2017) in SLATE and VQ-VAE (Van Den Oord et al., 2017) in SlotDiffusion, are employed to produce reconstruction targets for OCL training. Since these VAEs are designed for image generation, some methods adapt them for OCL. Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better. MSF (Zhao et al., 2024b) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs. Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
@@
Figure 2. VVO is a unified architecture that fully utilizes VFMs in OCL. It not only extracts VFM features with better objectness to facilitate object information aggregation; but further quantizes those VFM features as reconstruction targets to strengthen OCL training supervision. With typical SlotAttention or it variants as the aggregator and Vector-Quantization as the quantizer, VVO supports different VFMs as the encoder, and supports mainstream mixture, auto-regression and diffusion models as the decoder.
3. Proposed Method
@@
We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, or VQ-VFM-OCL (VVO), elegantly unifying mainstream OCL and consistently boosting their performance.
3.1. Unify OCL
@@
(4) $\bm{\phi}_{\mathrm{d}}:(\bm{S},\mathrm{destruct}(\bm{Q}))\rightarrow\bm{Q}^{\prime}$
Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (Kipf et al., 2022; Seitzer et al., 2023), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (Watters et al., 2019) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
Reconstructing $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible. Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important. We will explain this in Section 3.2.
@@
(5) $\mathrm{min}_{(\bm{\phi}_{\mathrm{a}},\bm{\phi}_{\mathrm{d}})}f_{\mathrm{recon}}(\bm{Q}^{\prime},\bm{Q})$
@@
3.2. Utilize VFMs in OCL
In our unified architecture, we utilize VFMs as following.
@@ 3.2. Utilize VFMs in OCL
We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly. No extra position encoding is needed here like in SA (Locatello et al., 2020) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$. Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly.
As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html. OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$. Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason. In Section 5.1, we analyze this both theoretically and empirically.
@@
(8) $\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$
(9) $\bm{I}=\mathrm{argmax}_{m}(\bm{P})$
(10) $\bm{Q}=\mathrm{index}_{m}(\bm{T},\bm{I})$
@@
For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ before Equation 8 to encourage code utilization:
(11) $\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature. We devise annealing residual connection (Zhao et al., 2024a, b) after Equation 10 to stabilize pre-training:
(12) $\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$
@@
(13) $l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
@@
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency and suppresses information redundancy or noise. This helps the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. After such tokenization, our reconstruction target can be either regressed using mixture and diffusion decoding or classified using auto-regressive decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision. In Section 5.2, we analyze this both theoretically and empirically.
3.3. Compare Architectures
Figure 3. Model architecture comparison.
As shown in Figure 3, we compare the baselines with our VVO in a unified perspective. Specifically,
- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2. (2) VFM features are shared quantized as reconstruction targets, which strengthens OCL supervision, as formulated in Section 3.2.
- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator. (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter. This loses much information of VFM features, while shared quantization can preserve such information (Van Den Oord et al., 2017).
- SLATE (the unofficially improved version; we adopt it) (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b) / STEVE (Singh et al., 2022c): (1) Same as our VVO. (2) Reconstruction targets are discretized from features of a separate VAE encoder, instead of quantized from features shared from the same VFM for OCL encoding, which causes optimization noises.
- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as our VVO. (2) Reconstruction targets are quantized from features of a separate VAE encoder, instead of sharing the same VFM for OCL encoding, which causes optimization noises.
- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as our VVO. (2) Reconstruction targets are shared from the same VFM for OCL encoding but without quantization, which causes optimization noises.
Our VVO adopts a clean and unified architecture with the above-mentioned two key designs. The first design, direct VFM feature extraction for OCL aggregation, is not proposed by us but we are the first to provide both theoretical and empirical analyses. The second design, shared VFM feature quantization as reconstruction targets, is proposed by us. Such shared quantization is also the key point that our VVO supports various OCL decoders. Specifically,
- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features. As quantized features are actually the clustering centroids of continuous features, these decoders are inherently compatible with our shared quantized VFM features.
- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are originally designed for discretized features, while (Seitzer et al., 2023) and (Kakogeorgiou et al., 2024) also explore their applicability to continuous features. Thus these decoders are naturally applicable to our shared quantized VFM features.
- Diffusion-based OCl decoders, e.g., Conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), are originally designed for dicretized features. Thus if without our shared quantized VFM features as reconstruction targets, then this type of OCL decoders would not be unified into our architecture.
Figure 4. Qualitative object discovery performance comparison.
resolution= ClevrTex #slot=11 COCO #slot=7 VOC #slot=6 MOVi-D #slot=21 conditional
256$\times$256 (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
SLATE-DINO 17.4
±2.9
87.4
±1.7
44.5
±2.2
43.3
±2.4
17.5
±0.6
28.8
±0.3
26.8
±0.3
25.4
±0.3
18.6
±0.1